darkknight:

1962, 75000) 1967, 120000) 1952, 55000) find the quadratic function using these points

darkknight:

I know how to do it. you use the ax^2+bx+c=y thing and plug each x value into the equation and set it equal to its corresponding y value. I know how to do 3 variable questions but I keep getting the wrong answer.

justjm:

Hm okay let me give it a try

darkknight:

k

darkknight:

the god

justjm:

The numbers they gave are like designed to trip anybody off bruh xD

darkknight:

I got random numbers but when I graphed them they weren't correct

darkknight:

I got c=1783402200 a=1400/3 b=-5473600/3

justjm:

Oh dear I got something else...I suppose it's getting late A=0.0101 B=1960.459 C=-3810299.942

justjm:

Well uh sorry I couldn't help on this one. The methodology is certainly correct and I used the same but the choice of numbers given in the question is quite simply horrendous.

darkknight:

I checked your answers but they don't work

aaa:

oof

Nnesha:

\[\rm 3849444a+1962b+C=75000 ~~~~~eq(1)\]\[\rm 3869089a+1967b+C=120000 ~~~~~eq(2)\]\[\rm 3810304a+1952b+C=55000 ~~~~~eq(3)\] subtract eq(1) &eq(2) will get eq(4) and eq(1) & eq(3) will get eq(5) \[\rm 19645a+5b=45000~~~~~eq(4)\] \[\rm 39140a+10b=20000~~~~~eq(5)\] solve eq(5) for b \[\rm b=\frac{ 20000-39140a }{ 10 }\] plugin in eq 4\[19645a+5(\frac{ 20000-39140a }{ 10 })=45000\] \[1945a+10000-19570a=45000 ~~~~\rightarrow 75a=45000-10000\]\[\rm a=466.667\] Substitute `a` values in \[\rm b=\frac{ 20000-39140a }{ 10 }\] and then a&b values in one of the 3 equations to get C value